# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ 2.2 Inferential objective and challenges in identification
where $\mathbb{E}_{0}^{\pi}$ denotes expectation under the counterfactual MDP induced by policy $\pi$. Throughout, we omit the superscript $\pi$ and write $q_{0}$ in place of $q_{0}^{\pi}$ when the policy under consideration is understood to be $\pi$. Intuitively, the $Q$-function represents the expected cumulative reward (or cost) for an individual who begins in state $s$, takes initial action $a$, and thereafter follows the policy $\pi$. The discount factor $\gamma$ determines the weight assigned to future rewards, controlling the time horizon for evaluation. By convention, we set $10^{0}:=$ 1; thus, when $\gamma=0$, the $Q$-function reduces to the short-term outcome regression $(a,s)\mapsto E_{0}[Y_{0}\mid A_{0}=a,S_{0}=s]$, and we recover the well-studied problem of inference on linear functionals of the outcome regression (Chernozhukov et al., 2018a, 2022; van der Laan et al., 2024b).
[…]
Statistically efficient estimation of the policy value, $E_{0}\left[V^{\pi}(q_{0})(S_{0})\right]$—a special case of a linear functional—via DRL has been studied under nonparametric models by van Der Laan et al. (2018), Kallus and Uehara (2020, 2022), and Tran et al. (2023). However, nonparametric inference on policy value is challenging, requiring strong conditions for identifiability and root-$n$ estimation. In particular, it depends on the existence and finite variance of the importance-weighted state occupancy ratio:
@@
where $\pi(a\mid s)$ is the target policy and $b_{0}(a\mid s)=P_{0}(A_{0}=a\mid S_{0}=s)$ is the behavior policy (Mehrabi and Wager, 2024). This condition requires overlap between the target and behavior policies, as well as between future and initial state distributions. Such overlap is often difficult to satisfy in high-dimensional settings, especially when interventions induce rare or novel states (D’Amour et al., 2021; Mehrabi and Wager, 2024), and even near-violations can cause high estimator variability. The asymptotic variance of existing DRL estimators and the Cramér–Rao bound depend critically on the variability of $d_{0}$.
A notable example of a linear functional of the $Q$-function is the long-term causal effect in a randomized experiment (Tran et al., 2023).
Example 2 (Long-term causal effect in A/B test).
Consider an A/B test where individuals are randomly assigned to one of two arms: either the intervention of interest or a control (such as no intervention or an alternative). In this case, the data-generating policy $b_{0}$ is static: $A_{t}:=Z$ for all $t\geq 0$, where $Z$ is the treatment arm indicator, with $Z=1$ representing the treatment arm and $Z=0$ the control. Following Tran et al. (2023), the long-term causal effect of treatment is defined as $\psi_{0}:=\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\gamma^{t}\{Y_{t}(1)-Y_{t}(0)\}\right],$ where $Y_{t}(1)$ and $Y_{t}(0)$ are the potential outcomes at time $t$ under treatment and control, respectively. Decompose the state as $S_{t}:=(Z,\widetilde{S}_{t})$, where $\widetilde{S}_{t}$ is a subvector of the state excluding the study assignment $Z$. Define the behavior policy as $b_{0}(a\mid(z,\widetilde{s})):=\mathbbm{1}\{a=z\}$. Then, $\psi_{0}$ can be expressed as $\psi_{0}=E_{0}\left[q_{0}^{b_{0}}(1,(1,\widetilde{S}_{0}))-q_{0}^{b_{0}}(0,(0,\widetilde{S}_{0}))\right],$ which is a linear functional of the $Q$-function $q_{0}^{b_{0}}$ corresponding to $m((z,\widetilde{s}),a,q)=q(1,(1,\widetilde{s}))-q(0,(0,\widetilde{s}))$. ∎
2.3 The Bellman equation as an integral equation
In this work, it is useful to consider an alternative identification of $q_{0}$ as the solution to a linear inverse problem. Let $\lambda$ be a measure on $\mathcal{S}\times\mathcal{A}$ that dominates the distributions of $(A_{0},S_{0})$ and $(A_{0},S_{1})$ almost surely for all $P\in\mathcal{P}$, and let $L^{\infty}(\lambda)$ denote the Banach space induced by the $\lambda$-essential supremum norm. By rearranging terms in (1), the $Q$-function $q_{P}$ for each $P\in\mathcal{P}$ is identified as the solution to the integral equation
[…]
where the outcome regression is defined by $\mu_{P}(a,s):=E_{P}[Y_{0}\mid A_{0}=a,S_{0}=s]$, and the Bellman integral operator $\mathcal{T}_{P}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ is defined as as $\mathcal{T}_{P}(h)(a,s):=h(a,s)-\gamma E_{P}\left[V_{h}^{\pi}(S_{1})\mid A_{0}=a,S_{0}=s\right].$ Consequently, our estimand $\psi_{0}$ corresponds to the $P_{0}$-evaluation of the target parameter $\Psi:P\mapsto E_{P}\left[m(S_{0},A_{0},q_{P})\right]$, defined over the nonparametric model $\mathcal{P}$. Inference on $\psi_{0}$ thus reduces to inference on a linear functional of the solution to a linear inverse problem (Ai and Chen, 2003, 2012; Bennett et al., 2022, 2023a, 2023b)—a perspective we exploit in our theoretical analysis.
@@ 3.1 Proposed estimator
In this section, we propose automatic DRL estimators for the estimand $\psi_{0}$ that impose semiparametric restrictions on the $Q$-function and outline their asymptotic properties. These estimators are automatic in the sense that they require only the specification of a linear functional, from which a debiasing procedure is derived. In later sections, we develop formal theory for semiparametric DRL and introduce model-robust and model-adaptive variants that remain valid under misspecification.
Suppose the $Q$-function $q_{0}$ lies in a (possibly infinite-dimensional) subspace $H\subset L^{\infty}(\lambda)$ of $\lambda$-essentially bounded functions, and let $\mathcal{P}_{H}:=\{P\in\mathcal{P}:q_{P}\in H\}$ denote the corresponding semiparametric model for $P_{0}$. For example, $q_{0}$ may be partially linear, additive, or depend only on a subset of the state variables. A possible choice for $H$ is the partially linear model (Robinson, 1988), which assumes $q_{0}(A_{0},S_{0})=q_{0}(0,S_{0})+A_{0}\beta_{0}^{T}S_{0}$. Here, the control function $q_{0}(0,S_{0})$ is estimated nonparametrically, while the treatment effect $q_{0}(1,S_{0})-q_{0}(0,S_{0})$ is modeled linearly in $S_{0}$. A natural estimator of $\psi_{0}=\Psi(P_{0})$ is the plug-in estimator $n^{-1}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})$, where $q_{n,H}\in H$ is an estimator of $q_{0}$. One approach to estimating $q_{0}$ is through fitted Q-iteration (FQI) (Munos and Szepesvári, 2008), which we discuss in Section 4. However, when $q_{n,H}$ is obtained via flexible learning methods, the plug-in estimator typically lacks $\sqrt{n}$-consistency and asymptotic normality due to first-order bias from $q_{n,H}-q_{0}$ (Van der Laan et al., 2011; Chernozhukov et al., 2018a). To address this, debiasing techniques are used to eliminate first-order bias and restore asymptotic properties.
To construct a bias correction, we require the linear functional $\psi_{0}=E_{0}[m(S_{0},A_{0},q_{0})]$ to be continuous not only in the $Q$-function $q_{0}$, but also in the outcome regression $\mu_{0}$. These quantities are linked through the linear inverse problem $\mathcal{T}_{0}(q_{0})=\mu_{0}$ in (3). To formalize this, we introduce the following notation and condition. For $P\in\mathcal{P}$, let $\langle\cdot,\cdot\rangle_{P}$ denote the $L^{2}(P_{A_{0},S_{0}})$ inner product, where $P_{A_{0},S_{0}}$ is the marginal distribution of $(A_{0},S_{0})$ under $P$, and let $\|\cdot\|_{P}$ be the associated norm. The closure of $H$ with respect to $\|\cdot\|_{P}$ is denoted $\overline{H}_{P}$, and we define $\mathcal{T}_{P}(\overline{H}_{P}):=\{\mathcal{T}_{P}(h):h\in\overline{H}_{P}\}$.
@@
As a consequence of Condition C1, we have $q_{0}=\mathcal{T}_{0}^{-1}(\mu_{0})$ and $\psi_{0}=E_{0}[m(S_{0},A_{0},\mathcal{T}_{0}^{-1}(\mu_{0}))]$, where the map $\mu\mapsto E_{0}[m(S_{0},A_{0},\mathcal{T}_{0}^{-1}(\mu))]$ is continuous on $\mathcal{T}_{0}(\overline{H}_{P_{0}})$. This condition implies that the functional $q\mapsto E_{0}[m(S_{0},A_{0},q)]$ is continuous with respect to the weak norm $\|\mathcal{T}_{0}(\cdot)\|_{P_{0}}$ induced by the Bellman operator, in the sense that
$\sup_{q\in H}\frac{E_{0}[m(S_{0},A_{0},q)]}{\|\mathcal{T}_{0}(q)\|_{P_{0}}}<\infty.$ (4)
In ill-posed inverse problems, conditions akin to (4) are often assumed directly to enable debiased estimation without requiring invertibility of $\mathcal{T}_{0}$ (Bennett et al., 2022). In our setting, however, the inverse problem defining $q_{0}$ is well-posed, and C1 holds under mild regularity conditions (Chen and Qi, 2022). Specifically, we show in Appendix A that $\mathcal{T}_{0}$ is a Fredholm operator of index zero, and that the inverse problem defining $q_{0}$ is a Fredholm equation of the second kind (Conway, 1994).
@@
The boundedness property in (4) yields a dual representation of the target parameter $\psi_{0}$ as a weighted expectation of the outcome $Y_{0}$ (Bennett et al., 2022, 2023b). The Riesz representation theorem ensures the existence of a representer $\alpha_{0,H}\in\overline{H}_{P_{0}}$ such that
$\psi_{0}=\langle\mathcal{T}_{0}(\alpha_{0,H}),\mathcal{T}_{0}(q_{0,H})\rangle_{P_{0}}=\langle\mathcal{T}_{0}(\alpha_{0,H}),\mu_{0}\rangle_{P_{0}},$
and hence $\psi_{0}=E_{0}[\mathcal{T}_{0}(\alpha_{0,H})(A_{0},S_{0})Y_{0}]$. The weighting function $\mathcal{T}_{0}(\alpha_{0,H})$ equals the inverse propensity weight function for the counterfactual mean $E_{0}[\mu_{0}(1,S_{0})]$ when $\gamma=0$, and equals the state occupancy ratio $d_{0}$ in (2) for the policy value estimand more generally. The Riesz representer is characterized as the minimizer of
$\alpha_{0,H}\in\arg\min_{\alpha\in\overline{H}_{P_{0}}}E_{0}\left[\left\{\mathcal{T}_{0}(\alpha)(A_{0},S_{0})\right\}^{2}-2m(S_{0},A_{0},\alpha)\right],$ (5)
a key fact underlying automatic Debiased Machine Learning (DML) (Chernozhukov et al., 2022; Bennett et al., 2023b; van der Laan et al., 2025).
Building on the weighted outcome representation, we propose an automatic DRL estimator that augments the plug-in estimator with a bias correction term. Specifically, given estimators $\widehat{\mathcal{T}}_{n}$, $\alpha_{n,H}$, and $q_{n,H}$ of $\mathcal{T}_{0}$, $\alpha_{0,H}$, and $q_{0}$, respectively, the autoDRL estimator $\psi_{n,H}$ takes the form of a one-step debiased estimator:
$\psi_{n,H}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H})(S_{1,i})-q_{n,H}(A_{0,i},S_{0,i})\right\},$ (6)
where the second term is as an influence-function-based bias correction. This estimator is automatic in the sense that its form is agnostic to the specific linear functional, requiring only an estimate of the Riesz representer, which can be learned using (5). Note that we do not require an estimator $\widehat{\mathcal{T}}_{n}$ of the entire operator $\mathcal{T}_{0}$, but only an estimate of the evaluation $\mathcal{T}_{0}(\alpha_{n,H})$. Estimation of the nuisance functions is discussed in Section 4. This estimator generalizes the nonparametric DRL estimator of Kallus and Uehara (2020) from policy value estimation to generic continuous linear functionals under semiparametric model restrictions. To accommodate the use of flexible machine learning methods, we recommend cross-fitting the nuisance estimators $q_{n,H}$ and $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$, following standard practice in debiased machine learning (van der Laan et al., 2011; Chernozhukov et al., 2018a). For simplicity, we omit the cross-fitting notation and refer the reader to these references for implementation details.
3.2 Asymptotic properties and efficiency considerations
We now study the asymptotic behavior of the DRL estimator $\psi_{n,H}$. We first show that $\Psi:P\mapsto E_{P}[m(S_{0},A_{0},q_{P})]$ is pathwise differentiable with respect to the semiparametric model $\mathcal{P}_{H}$ (Bickel et al., 1993). We then establish the asymptotic linearity of $\psi_{n,H}$ and discuss how leveraging semiparametric structure can yield efficiency gains over fully nonparametric estimators.
The following theorem establishes the pathwise differentiability of the parameter $\Psi$. For each $P\in\mathcal{P}_{H}$, we define the influence function

### v4

@@ 2.2 Inferential objective
where $\mathbb{E}_{0}^{\pi}$ denotes expectation under the counterfactual MDP induced by policy $\pi$ (kaelbling1996reinforcement). Throughout, when the policy is $\pi$, we omit the superscript and write $q_{0}$. The $Q$-function represents the expected cumulative reward for an individual starting in state $s$, taking action $a$, and subsequently following policy $\pi$. The discount factor $\gamma$ controls the weight placed on future rewards. Under the convention $0^{0}=1$, setting $\gamma=0$ reduces $q_{0}(a,s)$ to the reward function $r_{0}(a,s):=E_{0}[Y_{0}\mid A_{0}=a,S_{0}=s]$.
Our estimand is $\psi_{0}:=E_{0}[m(S_{0},A_{0},q_{0})]$, where $q\mapsto m(S_{0},A_{0},q)$ is a linear map. A key example is the long-term policy value $\mathbb{E}_{0}^{\pi}\!\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}(\pi)\right],$ which corresponds to $m(s,a,q)=\int q(a^{\prime},s)\,\pi(a^{\prime}\mid s)\,\nu(da^{\prime})$. The $Q$-function $q_{0}$ is identified from the observed data distribution $P_{0}$ as the fixed point of the Bellman equation (bellman1966dynamic; sutton1998reinforcement):
$q_{0}(A_{0},S_{0})=E_{0}\!\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid A_{0},S_{0}\right],$ (1)
where $V^{\pi}(q)(s^{\prime}):=\int q(a^{\prime},s^{\prime})\,\pi(a^{\prime}\mid s^{\prime})\,\nu(da^{\prime})$ denotes the value function. The Bellman equation expresses $q_{0}(A_{0},S_{0})$ as the sum of the immediate reward and the discounted future value $V^{\pi}(q_{0})(S_{1})$ under policy $\pi$. Setting $\gamma=0$ recovers the well-studied problem of inference on linear functionals of the outcome regression (chernozhukov2018double; chernozhukov2022automatic). A notable example of such a functional is the long-term causal effect in a randomized experiment (tran2023inferring).
Example 2 (Long-term causal effect in an A/B test).
Consider an A/B test in which individuals are randomly assigned to treatment or control, so the behavior policy is static: $A_{t}:=Z$ for all $t\geq 0$, where $Z\in\{0,1\}$ is the treatment indicator. Following tran2023inferring, the long-term causal effect is defined as $\psi_{0}:=\mathbb{E}_{0}[\sum_{t=0}^{\infty}\gamma^{t}\{Y_{t}(1)-Y_{t}(0)\}]$, where $Y_{t}(1)$ and $Y_{t}(0)$ denote the potential outcomes under treatment and control. Equivalently, $\psi_{0}=\psi_{0,1}-\psi_{0,0}$, with $\psi_{0,z}:=\mathbb{E}_{0}[q_{0}^{\pi_{z}}(z,S_{0})]$ for the deterministic policy $\pi_{z}(a\mid s)=\mathbbm{1}\{a=z\}$. Each component $\psi_{0,z}$ corresponds to the linear functional $m(s,a,q)=q(z,s)$. ∎
Challenges in nonparametric inference. Statistically efficient estimation of the policy value $E_{0}[V^{\pi}(q_{0})(S_{0})]$—a special case of a linear functional—via DRL has been studied under nonparametric models by van2018online, kallus2020double; kallus2022efficiently, and tran2023inferring. However, nonparametric inference is challenging because root-$n$ regular estimation requires the existence and finite variance of the unnormalized state–action discounted occupancy ratio
@@
where $b_{0}(a\mid s):=P_{0}(A_{0}=a\mid S_{0}=s)$ is the behavior policy (mehrabi2024off). This condition requires both policy overlap and intertemporal overlap, which can be restrictive in high-dimensional settings or when interventions induce rare or novel states (jin2021pessimism; d2021overlap; mehrabi2024off). Near-violations inflate estimator variance because the efficiency bound is determined by the variability of $d_{0}$ (kallus2022efficiently).
2.3 The Bellman equation as an inverse problem
We consider an equivalent representation of $\psi_{0}$ as a linear functional of the solution to a linear inverse problem (ai2003efficient; ai2012semiparametric; bennett2022inference; bennett2023minimax; bennett2023source). Let $\lambda$ be a measure on $\mathcal{S}\times\mathcal{A}$ that dominates the distributions of $(A_{0},S_{0})$ and $(A_{1},S_{1})$ for all $P\in\mathcal{P}$, and let $L^{\infty}(\lambda)$ denote the Banach space induced by the $\lambda$-essential supremum norm. By rearranging terms in (1), the $Q$-function $q_{P}$ for each $P\in\mathcal{P}$ is identified as the solution to the integral equation
$\mathcal{T}_{P}(q_{P})=r_{P},\quad\text{$P$-almost everywhere},$ (3)
where the reward function (or outcome regression) is defined by $r_{P}(a,s):=E_{P}[Y_{0}\mid A_{0}=a,S_{0}=s]$, and the Bellman integral operator $\mathcal{T}_{P}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ is defined as as $\mathcal{T}_{P}(h)(a,s):=h(a,s)-\gamma E_{P}\left[V^{\pi}(h)(S_{1})\mid A_{0}=a,S_{0}=s\right].$ It holds that $\psi_{0}=\Psi(P_{0})$ where we define target parameter $\Psi:P\mapsto E_{P}\left[m(S_{0},A_{0},q_{P})\right]$, over the nonparametric model $\mathcal{P}$.
@@ 3.1 Proposed estimator
In this section, we propose automatic DRL estimators for the estimand $\psi_{0}$ that impose semiparametric restrictions on the $Q$-function. In later sections, we describe how to modify these estimators to obtain model-robust inference and derive their asymptotic theory.
Suppose the true $Q$-function $q_{0}$ lies in a (possibly infinite-dimensional) subspace $H\subset L^{\infty}(\lambda)$, and define the corresponding semiparametric model $\mathcal{P}_{H}:=\{P\in\mathcal{P}:q_{P}\in H\}$. For example, $H$ may encode a partially linear structure of the form $q_{0}(A_{0},S_{0})=q_{0}(0,S_{0})+A_{0}\beta_{0}^{\top}S_{0}$, where the baseline function $q_{0}(0,S_{0})$ is estimated nonparametrically while the treatment contrast $q_{0}(1,S_{0})-q_{0}(0,S_{0})$ is modeled linearly in $S_{0}$ (robinson1988root). A natural estimator of $\psi_{0}=\Psi(P_{0})$ is the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})$, where $q_{n,H}\in H$ is an estimator of $q_{0}$. One way to obtain $q_{n,H}$ is through fitted $Q$-iteration, discussed in Section 3.5. However, when $q_{n,H}$ is estimated flexibly, the plug-in estimator generally fails to achieve $\sqrt{n}$-consistency or asymptotic normality because the first-order bias induced by $q_{n,H}-q_{0}$ dominates (van2011targeted; chernozhukov2018double). Debiasing is therefore required to eliminate this first-order bias and recover valid asymptotic inference.
To construct a bias correction, we require that the linear functional $\psi_{0}=E_{0}[m(S_{0},A_{0},q_{0})]$ be continuous in $q_{0}$, and that $q_{0}$ depend continuously on $r_{0}$ through the inverse relation $\mathcal{T}_{0}(q_{0})=r_{0}$. To formalize this, for $P\in\mathcal{P}$, let $\langle\cdot,\cdot\rangle_{P}$ denote the $L^{2}(P_{A_{0},S_{0}})$ inner product—where $P_{A_{0},S_{0}}$ is the marginal distribution of $(A_{0},S_{0})$ under PP—and let $\|\cdot\|_{P}$ be the associated norm. The closure of $H$ with respect to $\|\cdot\|_{P}$ is denoted $\overline{H}_{P}$, and we define $\mathcal{T}_{P}(\overline{H}_{P}):=\{\mathcal{T}_{P}(h):h\in\overline{H}_{P}\}.$ We impose the following conditions.
@@
Condition C1 ensures that $\psi_{0}=E_{0}[m(S_{0},A_{0},\mathcal{T}_{0}^{-1}(r_{0}))]$ is an $L^{2}$-continuous functional of the reward function $r_{0}$ on the range $\mathcal{T}_{0}(\overline{H}_{0})$, a standard requirement for regular estimation (chernozhukov2018learning; ichimura2022influence). Condition C1(ii) is standard in reinforcement learning and follows from the well-posedness of the Bellman equation (chen2022well). Appendix A further shows that $\mathcal{T}_{0}$ is a Fredholm operator of index zero and therefore admits a bounded inverse for all but finitely many discount factors $\gamma\in[0,1)$ (conway1994course).
@@
Condition C1 also implies that the functional $q\mapsto E_{0}[m(S_{0},A_{0},q)]$ is continuous with respect to the weak norm $\|\mathcal{T}_{0}(\cdot)\|_{P_{0}}$ induced by the Bellman operator, in the sense that
$\sup_{q\in H}\frac{|E_{0}[m(S_{0},A_{0},q)]|}{\|\mathcal{T}_{0}(q)\|_{P_{0}}}<\infty.$ (4)
@@
As a consequence, the target parameter $\psi_{0}$ admits a dual representation as a (signed) weighted expectation of the outcome $Y_{0}$. Specifically, by the Riesz representation theorem, there exists a representer $\alpha_{0,H}\in\overline{H}_{P_{0}}$ such that
$\psi_{0}=\langle\mathcal{T}_{0}(\alpha_{0,H}),\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}=\langle\mathcal{T}_{0}(\alpha_{0,H}),r_{0}\rangle_{P_{0}},$
and hence $\psi_{0}=E_{0}[\mathcal{T}_{0}(\alpha_{0,H})(A_{0},S_{0})Y_{0}]$. The weighting function $\mathcal{T}_{0}(\alpha_{0,H})$ generalizes the state–action occupancy ratio $d_{0}$ in (2) and plays a central role in debiasing.
We construct an automatic DRL estimator by augmenting the plug-in estimator with an influence-function-based bias correction (bickel1993efficient). Given estimators $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$ and $q_{n,H}$ of $\mathcal{T}_{0}(\alpha_{0,H})$ and $q_{0}$, respectively, we define the doubly robust estimator
$\psi_{n,H}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H})(S_{1,i})-q_{n,H}(A_{0,i},S_{0,i})\right\}.$ (5)
For the special case of the policy value and $H:=L^{\infty}(\lambda)$, this reduces to the nonparametric DRL estimator of kallus2020double. As in standard debiased machine learning, we recommend cross-fitting for nuisance estimation to mitigate overfitting and avoid Donsker conditions (van2011cross; chernozhukov2018double). For readability, we omit cross-fitting notation throughout.
The estimator (5) is automatic: its form does not depend on the choice of linear functional, but only on the Riesz representer. The representer optimizes the objective:
$\alpha_{0,H}=\arg\min_{\alpha\in\overline{H}_{P_{0}}}E_{0}\!\left[\{\mathcal{T}_{0}(\alpha)(A_{0},S_{0})\}^{2}-2\,m(S_{0},A_{0},\alpha)\right],$ (6)
which can be rewritten as a min–max saddle-point problem that avoids computing the conditional expectation inside $\mathcal{T}_{0}$ (see Section 3.5). Once $\alpha_{0,H}$ is estimated, the quantity $\mathcal{T}_{0}(\alpha_{0,H})$ is obtained via a second-stage regression. As in autoDML (chernozhukov2022automatic; bennett2023source; van2025automatic), the Riesz representer—and therefore the DRL estimator—can be learned using flexible machine learning methods for any choice of $m$.
3.2 Model-robust extension and asymptotic theory
Under semiparametric restrictions, the true $Q$-function $q_{0}$ may not lie exactly in the model $H\subset L^{\infty}(\lambda)$, but it may be well approximated by elements of $H$. Allowing small misspecification in exchange for efficiency gains can be advantageous, especially under limited overlap (imbensOverlapEstimand2006; van2023adaptive). For instance, linear MDP models yield substantial efficiency improvements (wang2020statistical; xie2023semiparametrically; wu2024statistical). More generally, $H$ may be adopted purely as a working model, with interest focused on the best approximation to $q_{0}$, such as the best linear predictor when $H$ is taken to be a linear class (whitney2020comment; vansteelandt2022assumption; chambaz2012estimation; chernozhukov2018generic).
To accommodate such settings, we extend our framework to model-robust inference on the projection estimand $\psi_{0,H}:=E_{0}\!\left[m(S_{0},A_{0},q_{0,H})\right],$ where $q_{0,H}$ is the Bellman projection of $q_{0}$ onto $H$. For any $P\in\mathcal{P}$, we define the Bellman projection
$q_{P,H}:=\arg\min_{q\in\overline{H}_{P}}E_{P}\!\left[\{Y_{0}-\mathcal{T}_{P}(q)(A_{0},S_{0})\}^{2}\right],$ (7)
that is, the best approximation to $q_{P}$ within $H$ under the norm $\|\mathcal{T}_{P}(\cdot)\|_{P}$. This induces a working parameter $\Psi_{H}(P):=E_{P}\bigl[m(S_{0},A_{0},q_{P,H})\bigr],$ which coincides with the original parameter $\Psi$ on $\mathcal{P}_{H}$ but remains well-defined over the full nonparametric model $\mathcal{P}$. The projected function $q_{P,H}$ satisfies the projected Bellman equation $\mathcal{T}_{P}(q_{P,H})=r_{P,H}$, where $r_{P,H}:=\arg\min_{r\in\mathcal{T}_{P}(\overline{H}_{P})}\|r_{P}-r\|_{P}$ is the $L^{2}(P)$-projection of the reward function $r_{P}$ onto the range $\mathcal{T}_{P}(\overline{H}_{P})$.
A model-robust and efficient estimator of the projection parameter $\Psi_{H}(P_{0})$ is obtained by adding a correction term to $\psi_{n,H}$:
$\psi_{n,H}^{*}$ $=\psi_{n,H}+\frac{1}{n}\sum_{i=1}^{n}\big(r_{n}-\widehat{\mathcal{T}}_{n}(q_{n,H})\big)(A_{0,i},S_{0,i})\left\{\alpha_{n,H}(A_{0,i},S_{0,i})-\gamma V^{\pi}(\alpha_{n,H})(S_{1,i})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\right\}.$
Here $q_{n,H}\in H$, $\alpha_{n,H}\in H$, $\widehat{\mathcal{T}}_{n}(q_{n,H})$, $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$, and $r_{n}$ are estimators of $q_{0,H}$, $\alpha_{0,H}$, $\mathcal{T}_{0}(q_{0,H})$, $\mathcal{T}_{0}(\alpha_{0,H})$, and $r_{0}$, respectively. The estimator $\psi_{n,H}^{*}$ generalizes the DRL estimator $\psi_{n,H}$ from the previous section, which is recovered under correct specification by setting $r_{n}:=\widehat{\mathcal{T}}_{n}(q_{n,H})$, in which case the additional correction term vanishes.
The asymptotic properties of $\psi_{n,H}^{*}$ (and of $\psi_{n,H}$ as a special case) follow from the next theorem, which establishes pathwise differentiability of $\Psi_{H}$, identifies its efficient influence function (EIF), and provides a von Mises expansion (mises1947asymptotic; bickel1993efficient). For each $P\in\mathcal{P}$, we define the function
$\varphi^{*}_{P,H}(s,a,y,s^{\prime})$ $:=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\,\{y+\gamma V^{\pi}(q_{P,H})(s^{\prime})-q_{P,H}(a,s)\}$
$\qquad+\bigl\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(s^{\prime})-\mathcal{T}_{P}(\alpha_{P,H})(a,s)\bigr\}\bigl\{r_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\bigr\}$
$\qquad+m(s,a,q_{P,H})-\Psi_{H}(P).$